Mr. K’s is a very popular hair salon. It offers high-quality hairstyling and physical relaxation services at a reasonable price, so it always has unlimited demand. The service process includes five activities that are conducted in the sequence described next (the time required for each activity is shown in parentheses):

  • Activity 1: Welcome a guest and offer homemade herb tea (11 minutes).
  • Activity 2: Wash and condition hair (11 minutes).
  • Activity 3: Neck, shoulder, and back stress-release massage (11 minutes).
  • Activity 4: Design the hairstyle and do the hair (24 minutes).
  • Activity 5: Check out the guest (7 minutes).

Three servers (S1, S2, and S3) offer the services in a worker-paced line. The assignment of tasks to servers is the following: S1 does Activity 1, S2 does activities 2 and 3, and S3 does activities 4 and 5.

Mr. K’s is a very popular hair salon. It offers high-quality hairstyling and physical relaxation services at a reasonable price, so it always has unlimited demand. The service process includes five activities that are conducted in the sequence described next (the time required for each activity is shown in parentheses):

  • Activity 1: Welcome a guest and offer homemade herb tea (11 minutes).
  • Activity 2: Wash and condition hair (11 minutes).
  • Activity 3: Neck, shoulder, and back stress-release massage (11 minutes).
  • Activity 4: Design the hairstyle and do the hair (24 minutes).
  • Activity 5: Check out the guest (7 minutes).

Three servers (S1, S2, and S3) offer the services in a worker-paced line. The assignment of tasks to servers is the following: S1 does Activity 1, S2 does activities 2 and 3, and S3 does activities 4 and 5.

Mr. K’s is a very popular hair salon. It offers high-quality hairstyling and physical relaxation services at a reasonable price, so it always has unlimited demand. The service process includes five activities that are conducted in the sequence described next (the time required for each activity is shown in parentheses):

  • Activity 1: Welcome a guest and offer homemade herb tea (11 minutes).
  • Activity 2: Wash and condition hair (11 minutes).
  • Activity 3: Neck, shoulder, and back stress-release massage (11 minutes).
  • Activity 4: Design the hairstyle and do the hair (24 minutes).
  • Activity 5: Check out the guest (7 minutes).

Three servers (S1, S2, and S3) offer the services in a worker-paced line. The assignment of tasks to servers is the following: S1 does Activity 1, S2 does activities 2 and 3, and S3 does activities 4 and 5.

Mr. K’s is a very popular hair salon. It offers high-quality hairstyling and physical relaxation services at a reasonable price, so it always has unlimited demand. The service process includes five activities that are conducted in the sequence described next (the time required for each activity is shown in parentheses):

  • Activity 1: Welcome a guest and offer homemade herb tea (11 minutes).
  • Activity 2: Wash and condition hair (11 minutes).
  • Activity 3: Neck, shoulder, and back stress-release massage (11 minutes).
  • Activity 4: Design the hairstyle and do the hair (24 minutes).
  • Activity 5: Check out the guest (7 minutes).

Three servers (S1, S2, and S3) offer the services in a worker-paced line. The assignment of tasks to servers is the following: S1 does Activity 1, S2 does activities 2 and 3, and S3 does activities 4 and 5.

Mr. K’s is a very popular hair salon. It offers high-quality hairstyling and physical relaxation services at a reasonable price, so it always has unlimited demand. The service process includes five activities that are conducted in the sequence described next (the time required for each activity is shown in parentheses):

  • Activity 1: Welcome a guest and offer homemade herb tea (11 minutes).
  • Activity 2: Wash and condition hair (11 minutes).
  • Activity 3: Neck, shoulder, and back stress-release massage (11 minutes).
  • Activity 4: Design the hairstyle and do the hair (24 minutes).
  • Activity 5: Check out the guest (7 minutes).

Three servers (S1, S2, and S3) offer the services in a worker-paced line. The assignment of tasks to servers is the following: S1 does Activity 1, S2 does activities 2 and 3, and S3 does activities 4 and 5.

a. what is the labor content?

b. what is the average labor utilization?

c. at a wage rate of $21 per hour, what is the cost of direct labor per customer?

d. Mr. K considers hiring a new employee to help any one (and only one) of the servers without changing the tasks performed by each server. what will be the new direct labor?

e. Mr. K contemplates redesigning the assignment of tasks to servers. For this, Mr. K is evaluating the reassignment of Activity 5 from S3 to S1. what will be the new direct labor?

f. Mr. K is thinking to add one additional worker to the process. the worker would be assigned to the same set of tasks as one of the current workers. first, decide which set of tasks would benefit from one additional worker, then calculate the process capacity (customers per hour)?

g. suppose 4 minutes of processing time from activity 4 could be moved to activity 3 while keeping the task assignments the same. what would be the capacity of this new process (customers per hour)?

h. suppose each worker could be trained to complete all activities and they could all work independently of each other. what would be the capacity of this process (customers per hour)?
Answer #1

a. Total labor content = Sum of all activity time = 11+11+24+11+7 = 64 minutes

b) Total time S1 takes = 11 minutes

Total time S2 takes = 22 minutes

Total.time S3 takes = 31 minutes

Bottleneck = S3 resource

Now idle time at S1 = 31-11= 20 minutes

Idle time at S2 = 31- 22 = 9 minutes

Total idle time = 29 minutes / customer

Average labor utilization = ( Total labor content) / ( Total labor content + Idle time) = 64/ 93 =68.82%

C) Number of customer served per 60 minutes = 60 / bottleneck time = 60/31 = 1.94 customers

Total wage rate / hour = 21*3 = $63

Direct labor / customer = 63/1.94 = $32.47 ( note if you don't round off number of customers to two decimal places, the answer can change)

D) if one more resource is added, it will assist the bottleneck S3. The activity time at S3 will be halved and bottleneck shifts to resource S2.

New total number of customers served per hour = 60 / 22 = 2.73 customers

New direct labor cost per hour = 4*21 = $84

New direct labor cost / customer = 84 / 2.73 = $30.77 per customer

Note : solved first four parts as per Q-a policy.